OpenStudy (anonymous):

Factor. m2 + 12m + 36 (Points : 1) (m + 6)^2 (m – 6)^2 (m + 36)^2 Cannot be factored

OpenStudy (anonymous):

D , cannot be factored

OpenStudy (whpalmer4):

Not true!

OpenStudy (whpalmer4):

@dsfasdfasdfga please come back and I'll teach you how to factor this.

OpenStudy (whpalmer4):

@Briaa420 that invitation applies to you as well

OpenStudy (anonymous):

Alrighty im listening

OpenStudy (whpalmer4):

To factor something of the form \[ax^2+bx+c\]multiply the values of \(a\) and \(c\) together. For this problem, \(a =1, b = 12, c = 36\), so \(a*c = 36\) Now, find two factors of \(a*c\) that add up to \(b\). 36= 1*36 36 = 2*18 36 = 3*12 36 = 4*9 36= 6*6 6+6 = 12, so we will use 6 and 6. Rewrite the polynomial, splitting the middle term into two terms using those factors we found: \[m^2 + 6m + 6m + 36\]Now group each pair of terms: \[(m^2+6m) + (6m + 36)\]Factor each group separately:\[m(m+6) + 6(m+6)\]Look at that — each group has a common factor of \((m+6)\), so factor it out: \[(m+6)(m+6) = (m+6)^2\] Check by multiplying out the factoring to make sure we get our original back: \[(m+6)(m+6) = m(m+6) + 6(m+6) = m*m+6m + 6m + 36\]\[\qquad=m^2+12m+36\checkmark\]

OpenStudy (anonymous):

i never knew that ! so which one would be the answer choice?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

Oh its (m+6) ^2

OpenStudy (anonymous):

A

OpenStudy (whpalmer4):

Yes, A would be the correct answer choice: \[(m+6)^2 = (m+6)(m+6) = m(m+6) + 6(m+6) = m^2 + 6m + 6m + 36\]\[\qquad = m^2+12m+36\]